OpenStudy (anonymous):

Solve by completing the square. h2 – 22h = -29

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

what is half of 22?

OpenStudy (misty1212):

not a trick question, you need that first

OpenStudy (anonymous):

11 right

OpenStudy (misty1212):

yes and what is \(11^2\)?

OpenStudy (anonymous):

5.5

OpenStudy (misty1212):

oh no

OpenStudy (misty1212):

not half of 11, 11 squared, ie. \(11\times 11\)

OpenStudy (anonymous):

so is it h =(22+√608)/2=11+2√ 38 = 23.329

OpenStudy (misty1212):

i have no idea, we are not there yet

OpenStudy (misty1212):

looks like you used the quadratic formula instead of completing the square